mathslover (mathslover):

Mathematics Quiz #2 : Get ready. Total Questions :10 Marks for correct answer : 4 Marks for wrong answer : -1 The first who answers will be awarded with +4. The second who answers with a short hint of the soln will also be awarded +4. Only 2 answers will be considered. Are you guys ready? [part 2]

OpenStudy (anonymous):

c'mon!

OpenStudy (anonymous):

:-p

mathslover (mathslover):

Solve the equation : \((6-x)^4 + (8-x)^4 = 16\)

OpenStudy (raden):

use wolfram, lol

mathslover (mathslover):

Oh... Don't cheat urself and others PLEASE

OpenStudy (anonymous):

okay fine :-(

terenzreignz (terenzreignz):

LOL

OpenStudy (anonymous):

(6,0) (8,0)

mathslover (mathslover):

@chris00 , give me a hint...

OpenStudy (anonymous):

graph it and find the roots

mathslover (mathslover):

Not proper method. I guess u used wolfram or any other source?

OpenStudy (anonymous):

graph (6-x)^4+(8-x)^4-16=0

OpenStudy (anonymous):

my calculator

terenzreignz (terenzreignz):

okay, I just lost it.... LOL

OpenStudy (anonymous):

thats a proper source.

mathslover (mathslover):

lol Your answer will not be considered. [AS you used calculator] Yes @SharpEdge , changing the variable : +4

OpenStudy (anonymous):

rules should be more explicit. haha

mathslover (mathslover):

One more chance left. Oldrin, don't type whole soln, just give a hint...

mathslover (mathslover):

1 minute more...

mathslover (mathslover):

Time is up.

mathslover (mathslover):

Points table : @SharpEdge : 14 @oldrin.bataku : 14 @terenzreignz : 12 @johnweldon1993 : 4 @chris00 : 1

mathslover (mathslover):

Next question : 7 , be ready

OpenStudy (anonymous):

least i'm in the positives

terenzreignz (terenzreignz):

So much for four points all, you only gave @oldrin.bataku 2 points...

OpenStudy (anonymous):

It's good! Just post the questions faster I have a test tomorrow morning and it's almost 12 ;-)

mathslover (mathslover):

Okay.. Solve : \(x^2 - 5|x| + 6 = 0 \)

terenzreignz (terenzreignz):

+/- 1 +/- 6

OpenStudy (anonymous):

$$u=|x|\\u^2-5u+6=0\\u=3,2\\x=-3,-2,2,3$$

OpenStudy (raden):

x=2,3

mathslover (mathslover):

No^ (-1) @terenzreignz

terenzreignz (terenzreignz):

crud. :D me too much of in a hurry

mathslover (mathslover):

@oldrin.bataku : +4 @RadEn : (-1)

OpenStudy (anonymous):

that was mighty quick terenz.

terenzreignz (terenzreignz):

too quick. oh well :D

OpenStudy (raden):

ah, +-2, +-3 i mean :)

terenzreignz (terenzreignz):

New thread?

terenzreignz (terenzreignz):

I was TJ.... still am, really XD

OpenStudy (anonymous):

No need for a new thread yet... btw nice try @RadEn ;-) hehe

mathslover (mathslover):

Yes... @SharpEdge : +4 @RadEn [ U had to post the hint too :( ]

OpenStudy (raden):

if x > 0, then x = 2,3 if x < 0, then x = -2,-3

mathslover (mathslover):

Points table : @oldrin.bataku : 18 @SharpEdge : 18 @terenzreignz : 11 @johnweldon1993 : 4 @chris00 : 1 @RadEn : -1

mathslover (mathslover):

RadEn, sorry a bit late :(

OpenStudy (raden):

haha, nope :)

OpenStudy (karatechopper):

can I enter and try?

OpenStudy (anonymous):

entry fee is 5 dollars

terenzreignz (terenzreignz):

Better question would be who can stop you @karatechopper :D

mathslover (mathslover):

Yep KC. You're welcome. But read the rules first... :)

OpenStudy (karatechopper):

I give u a high five. @chris00

mathslover (mathslover):

Question no. 8 : Solve the equation : |x-|4-x||- 2x = 4

OpenStudy (anonymous):

that will do

mathslover (mathslover):

Suggestion : Don't be in hurry (@terenzreignz ) :P

OpenStudy (anonymous):

(0,0)

mathslover (mathslover):

Post the hint too. + Answer

OpenStudy (anonymous):

$$x=0$$

mathslover (mathslover):

@chris00 , did u mean 0 as the answer?

OpenStudy (anonymous):

yes

mathslover (mathslover):

@chris00 : + 4 @SharpEdge : -1 @oldrin.bataku [give hint, if u want +4] :)

OpenStudy (karatechopper):

I got 0

mathslover (mathslover):

KC hint please..

OpenStudy (anonymous):

$$|x-4+x|-2x=4\\2x-4-2x=4\\-4=4\\|x-x+4|-2x=4\\4-2x=4\\x=0$$

mathslover (mathslover):

Time up for all. @oldrin.bataku : +4 Points table : @oldrin.bataku : 22 @SharpEdge : 17 @terenzreignz : 11 @johnweldon1993 : 4 @chris00 : 5 @RadEn : -1

terenzreignz (terenzreignz):

conflicting answers... interesting :D

mathslover (mathslover):

@karatechopper , u have to give hint also... :(

OpenStudy (anonymous):

no conflicting answers here...:-)

OpenStudy (karatechopper):

ohh I get it now..

mathslover (mathslover):

Be ready for second last question...

terenzreignz (terenzreignz):

oh wait, sharpedge got -1 nvm

OpenStudy (karatechopper):

its k. Don't put me in points I'm just wanting to learn.

mathslover (mathslover):

Good spirit KC. Loved it

OpenStudy (anonymous):

hahahaha gd one.

terenzreignz (terenzreignz):

no, I thought you got it right, and hence, I thought we got conflicting answers

OpenStudy (raden):

ah, algebra again.. where is these : combinatoric or trigono/geo or others hahaha

terenzreignz (terenzreignz):

Possibly in the next few quizzes to come, if ever

OpenStudy (anonymous):

yes more!! do the last questions c'mon :-)

mathslover (mathslover):

How many number of 6 digits can be formed with the digits 1,2,3,4,5,6 in which 5 always occur in ten's place ? @RadEn : combinatorics :)

mathslover (mathslover):

This is the second last Sharp.

terenzreignz (terenzreignz):

6^5

terenzreignz (terenzreignz):

Principle of Multiplication?

mathslover (mathslover):

@SharpEdge : +4 @terenzreignz : Nopes.. (-1)

mathslover (mathslover):

One more chance left..

terenzreignz (terenzreignz):

You could have mentioned that no digits may be repeated

OpenStudy (anonymous):

$$5\ \_\ \_\ \_\ \_\ \_$$With \(5\) free spots each having one less digit we get \(5\times4\times3\times2\times1=120\)

mathslover (mathslover):

If u repeat any digit then one digit will be left. So no need to mention that.

OpenStudy (anonymous):

one less digit possibility than the previous

mathslover (mathslover):

@oldrin.bataku : +4

terenzreignz (terenzreignz):

You didn't say you had to use all the digits neither... c'mon @mathslover a little clarity next time?

mathslover (mathslover):

lol Okay... Sorry :( :)

terenzreignz (terenzreignz):

At the very least, retract that -1... (pretty please)

mathslover (mathslover):

lol

OpenStudy (anonymous):

assert urself terenz.

terenzreignz (terenzreignz):

I need ice cream.

OpenStudy (anonymous):

@terenzreignz has a point... I almost said \(6^5\) then I saw the other guy's answer accepted and knew immediately every digit had to be used

mathslover (mathslover):

@oldrin.bataku : 26 @SharpEdge : 21 @terenzreignz : 11 [Regreted.] @johnweldon1993 : 4 @chris00 : 5 @RadEn : -1